a thermometer has long calibration it reads the melting point as minus 10 degree it reads 60 degree in place of 50 degree Celsius calculate the temperature of boiling point of water in the 
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
(60-(-10))/(B.P -(-10))= (50-0)/(100-0)So, B.P =130°C. 
Regards
Arun (askIITians forum expert)